In the collection of nickels dimes, and quarters, there are twice as many dimes as nickels, and 2 fewer quarters than dimes. The total value of the coins is $4. How many nickels are there in the collection?
Answer by mananth(16949)

there are twice as many dimes as nickels, and 2 fewer quarters than dimes. The total value of the coins is $4. How many nickels are there in the collection?
let nickels be x
dimes = 2x
quarters = 2x-2
...
5x+10*2x+25(2x-2)=400
5x+20x+50x-50=400
75x=400+50
75x=450
x=6 nickels
dimes= 2x=12
quarters= 2x-2 = 10
..
6*5+12*10+10*25=400
...
m.ananth@hotmail.ca